solve: dx/(y^2 + z^2) = dy/-xy = dz/ -xz


1
Expert's answer
2021-05-07T14:17:06-0400

Given differential equation is-


"\dfrac {dx}{(y^2 + z^2)} = \dfrac{dy}{-xy} = \dfrac{dz}{ -xz}"


Taking Last two terms-


"\dfrac{dy}{-xy}=\dfrac{dz}{-xz}\\[9pt]\Rightarrow \dfrac{dy}{y}=\dfrac{dz}{z}"


Integrate the above equation and we get-


"lny=lnz+lnc_1\Rightarrow lny=lnc_1z\Rightarrow y=c_1z~~~~~-(1)"


Also "c_1=\dfrac{y}{z}~~~~~~-(2)"


Now Taking first and last terms-


"\dfrac{dx}{y^2+z^2}=\dfrac{dz}{-xz}\\[9pt]\Rightarrow \dfrac{ xdx}{z^2c_1^2+z^2}=\dfrac{dz}{-z}~~~~~~~~~~~~~~~~~~~~\text{ from eqn.(1)}\\[9pt]\Rightarrow xdx=-(c_1^2+1)zdz"


Integrating Both the sides-


"\dfrac{x^2}{2}=\dfrac{-z^2}{2}(c_1+1)^2+c_2"


"c_2=\dfrac{x^2}{2}+\dfrac{z^2(\dfrac{y}{z}+1)^2}{2}~~~~~~~~~~\text{from eqn.(2)}\\[9pt]=\dfrac{x^2}{2}+\dfrac{(y+z)^2}{2}~~~~~~~~~-(3)"


The solution of the equation is-


"\phi(c_1,c_2)=0
\\[9pt]


\phi(\dfrac{y}{z},\dfrac{(y+z)^2}{2})=0"